A horse pulls a wagon with a force of 50 N at an angle of 37o with the horizontal and moves along at a speed of 6 km/hr. (a) How much work does the horse do in 10 minutes? (b) What is the power output of the horse?


1
Expert's answer
2021-11-24T11:50:22-0500

(a) Work:


"W=Fvt\cos\theta=33.3\text{ kJ}."

(b) Power:

"P=Fv\cos\theta=66.6\text{ W}."